OpenStudy (anonymous):

how did the expression: (x-2)^3 - 8y^3 when expanded became: (x-2-2y)[(x^2-4x+4)+2(xy-2y)+4y^2]? Help please?

jimthompson5910 (jim_thompson5910):

they are using the formula \[\Large a^3 - b^3 = (a-b)(a^2 + ab + b^2)\] difference of cubes formula

jimthompson5910 (jim_thompson5910):

in this case a = x-2 b = 2y ... since (2y)^3 = 8y^3

OpenStudy (anonymous):

let me try it. :D

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

I GOT IT. thanks a lot! :)

jimthompson5910 (jim_thompson5910):

you're welcome